OpenStudy (mckenzieandjesus):

Simplify the difference (7w2 – 7w – 8) – (5w2 + 2w – 3) I got (w-5)(2w+1) but thats not on my answer choices.. 12w2 + 9w + 5 2w2 – 9w – 5 2w2 – 5w – 11 12w2 – 5w – 11

OpenStudy (mckenzieandjesus):

@arivillafana

OpenStudy (mckenzieandjesus):

@jim_thompson5910

OpenStudy (mckenzieandjesus):

@campbell_st

OpenStudy (jdoe0001):

\(\bf (7w^2-7w-8)-{\color{brown}{ (5w^2+2w-3) }} \\ \quad \\ (7w^2-7w-8){\color{brown}{ -5w^2-2w+3 }} \\ \quad \\ 7w^2-7w-8-5w^2-2w+3\implies ?\)

OpenStudy (mckenzieandjesus):

10w?

OpenStudy (jdoe0001):

10w? whatever happened to the other two terms?

OpenStudy (mckenzieandjesus):

huh?

OpenStudy (jdoe0001):

reduce the like terms notice the like terms \(\bf {\color{blue}{ 7w^2}}{\color{brown}{ -7w}}-8{\color{blue}{ -5w^2 }}{\color{brown}{ -2w }}+3\implies ?\)

OpenStudy (campbell_st):

just group the like terms \[7w^2 - 5w^2 = ?\] -7w - 2w = ? -8 + 3 and what's left is the solution

OpenStudy (mckenzieandjesus):

\[7w^2-5w^2=24 -7w-2w=-9 -8+3=-5\]

OpenStudy (mckenzieandjesus):

What would I do next?

OpenStudy (anonymous):

(7w^2 – 7w – 8) – (5w^2 + 2w – 3) 7w^2 - 7w -8 - 5w^2 - 2w + 3 Distributed the negative 7w^2 - 5w^2 - 7w - 2w - 8 + 3 Rearranged 2w^2 - 7w - 2w - 8 + 3 Solved 7w^2 - 5w^2 2w^2 - 9w - 8 + 3 Solved -7w - 2w 2w^2 - 9w - 5 <---Answer Solved -8 + 3

OpenStudy (mckenzieandjesus):

Oh yay thats what i have marked down already :)

OpenStudy (campbell_st):

\[7w^2 - 5w^2 = 2w^2\] 7w - 2w = 5w -8 + 3 = -5 so the answer is \[2w^2 + 5w - 5 \]